# Einstein-Horndeski gravity and the ultra slowly evaporating black hole

### v1

@@ 1 Introduction
Through indirect and direct observations by LIGO and the Event Horizon Telescope [1, 2, 3, 4], black holes, the most extraordinary objects in the universe, have received increasing attention. One of the most interesting things about black holes is that they behave like a thermodynamical system. In Ref.[5, 6], by using quantum field theory in curved space-time, Hawking showed that quantum effects allow black holes to radiate like black bodies which are now called Hawking radiation. From the Hawking radiation spectrum, one can calculate the power of the particle emission, investigate the evaporation behavior, and estimate the lifetime of black holes. After the discovery of Hawking radiation, Page gave numerical results for the lifetimes of uncharged and non-rotating black holes in asymptotically flat spacetime [7]. He found that there is a simple relation between a black hole’s lifetime and its initial mass $t\sim M_{0}^{3}$ [7], and then extended these discussions to rotating black hole case [8]. In Ref.[9], by combining the Hawking radiation and Schwinger effect in quantum electrodynamics, Hiscock and Weems generalized the discussions of black hole evaporation to the charged black hole case. They found that the presence of electric charge may greatly extend the lifetime of a black hole [9]. The study of the evaporation behavior of black holes has been a fruitful area up until now, it has been extended to many different spacetime backgrounds, such as de-Sitter spacetimes [10] and anti-de-Sitter spacetimes [11]. Due to the diversity of gravitational theories, there are also many discussions on black hole evaporation in modified gravity theories [12, 13, 14, 15]. Moreover, black hole evaporation may also be of help in the search for dark matter and dark energy [16].
@@
Einstein’s general relativity where all the matter fields are minimally coupled to gravity has passed various direct or indirect tests. However, many extensions of general relativity have also been taken into account in order to explain the accelerated expansion of our universe [17, 18]. Particularly, a lot of work has been devoted to finding the most general non-minimal coupled scalar-tensor action whose equations of motion have at most second-order derivatives. In the 1970s, general classes of theories of this kind were constructed by Horndeski, which is now called Einstein-Horndeski gravity [19]. Although Horndeski gravity is a scalar-tensor theory, by gauging the shift symmetry of the scalar field, it is also equivalent to the Einstein-vector theory which is shown in Ref.[20]. Horndeski gravity has the coupling between the derivative of the scalar field and Einstein tensor. This feature makes Horndeski gravity theory useful in inflationary cosmology [21, 22] and late-time cosmology [23]. Besides investigations in the cosmological background, Horndeski gravity has also received much attention in static and spherically symmetric black hole systems. Various interesting topics have been discussed in Horndeski black hole background. To name a few, Ref.[24, 25, 26] discussed the Horndeski black hole thermodynamics, Ref.[27, 28, 29] discussed the holographic implication of Horndeski black hole, and [30] discussed the black hole shadow in Horndeski background. Despite the above progress, black hole evaporation behavior which is an important part of black hole thermodynamics has not been investigated in Horndeski gravity background.
@@
In this work, we only consider the static and spherically symmetric black hole solution. So we can write the most general ansatz as
@@
in which the $\Phi(r)$ is the U(1) gauge field potential. After this simplification, there are three independent equations from the above equations (2)
@@
$4(2rf^{\prime}+2(f-1)+2\Lambda r^{2})+\frac{2q^{2}}{r^{2}}+2\alpha r^{2}f\phi^{\prime 2}$
$+2\gamma(4rf\phi^{\prime\prime}+(3rf^{\prime}+(f+1))\phi^{\prime})f\phi^{\prime}$ $=0,$
$4(2rfh^{\prime}+2h(f-1)+2\Lambda r^{2}h)+\frac{2q^{2}}{r^{2}}h-2\alpha r^{2}fh\phi^{\prime 2}$
@@
As we are mostly interested in asymptotically Minkowski black hole solution, we will focus on the case where $\alpha=0$ and $\Lambda=0$. In this scenario, an asymptotically flat black hole solution can be obtained where [31, 26]
@@
$h(x)=1-\frac{2}{x}+\frac{4y^{2}}{x^{2}}-\frac{4y^{4}}{3x^{4}}$ (10)
The charge parameter $y$ of this Horndeski black hole always satisfies the constraint Eq.(8). Under this constraint, we plot the $h(r)$ in Fig.1 to show the black hole metric function for different values of charge parameter $y$. It is easy to find that the metric function has only one line intersecting the $z=0$ plane which means that there is only one positive real root of Eq.(9). Thus the Horndeski black hole has only one event horizon which is different from the charged black hole in Einstein gravity.
@@ 3 Thermodynamics of charged black holes in Einstein-Horndeski gravity
$\int_{c}\omega(\tilde{\phi},\delta\tilde{\phi},\mathcal{L}_{\xi}\tilde{\phi})=\delta\mathcal{H}_{\infty}-\delta\mathcal{H}_{+}$ (12)
@@
$\delta\mathcal{H}_{\infty}=\delta M-\Phi_{e}\delta Q$ (13)
@@
$\delta\mathcal{H}_{+}=T\delta S+\Phi_{\phi}\delta Q_{\phi}$ (16)
Due to this additional term, the thermodynamic first law is different from the RN case which reads
@@
$\Phi_{e}=\frac{4Q}{{r_{h}}}-\frac{8Q^{3}}{3r_{h}^{3}}.$
The condition $\eta<0$ is required to ensure these thermodynamic variables are real. Moreover, unlike Einstein gravity, the entropy of the black hole in Horndeski gravity has correction which is
@@ 4 Evaporation behavior of charged black holes in Einstein-Horndeski gravity:
$(\alpha g^{rr}-\eta G^{rr})\phi^{\prime}\equiv 0$ (21)
which means the $\phi$ field can not be solved by its equation of motion. This is very different from other cases where scalar hair is involved such as Ref.[32, 33, 34]. In these cases, the scalar field is solved by its equation of motion.
@@
$\frac{dM}{dt}=T\frac{dS}{dt}+\Phi_{e}\frac{dQ}{dt}+\Phi_{\phi}\frac{dQ_{\phi}}{dt}$ (25)
@@
The first term regarding thermal radiation follows Stefan-Boltzmann law.
$\frac{\delta Q_{H}}{dt}=T\frac{dS}{dt}=-\sigma T^{4}$ (26)
@@
$\frac{dr}{d\lambda}=\sqrt{\frac{f(r)}{b^{2}h(r)}-\frac{f(r)}{r^{2}}}.$ (29)
where $b=L/E$ is called the impact factor. A massless particle that can reach infinity must obey the equation
@@
$\frac{\mathrm{d}M}{\mathrm{d}t}=-\frac{\pi^{3}}{15}b_{c}^{2}T^{4}+\Phi_{e}\frac{\mathrm{d}Q}{\mathrm{d}t}+\Phi_{\phi}\frac{\mathrm{d}Q_{\phi}}{\mathrm{d}t}.$ (34)
@@
$\frac{\mathrm{d}Q_{\phi}}{\mathrm{d}t}=\frac{\partial Q_{\phi}}{\partial M}\frac{\mathrm{d}M}{\mathrm{d}t}+\frac{\partial Q_{\phi}}{\partial Q}\frac{\mathrm{d}Q}{\mathrm{d}t}$ (35)
@@
To understand the difference in evaporation behavior between RN and Horndeski black hole, we see that in combination with (37) and (39), the charge loss rate is closely related to the electric field strength outside the event horizon. At the early stage of evaporation, the horizon radius $r_{h}$ is large, thus $O(1/r^{4})$ term can be neglected in Eq.(38). Therefore, the electric field strength of Horndeski black holes outside the event horizon is the same as the RN case which is both $O(\frac{1}{r^{2}})$ Coulomb type. As the black hole evaporates, $r_{h}$ becomes smaller and the $O(\frac{1}{r^{4}})$ term of Horndeski electric field strength (38) becomes more and more important. As can be seen from Fig.2, at the late stage of evaporation, the exterior electric field strength of the Horndeski black hole is vastly different from the RN case. The electric field strength outside Horndeski black hole first increases and then decreases instead of being a monotonically decreasing function. This difference in the electric field strength is the crucial reason for why the Horndeski black holes have a much longer lifetime than RN black holes.
@@ 5 Conclusion
For future directions, firstly, we only consider the asymptotically flat Horndeski black hole in this work. For Horndeski gravity, there are also black hole solutions in asymptotically AdS spacetime [26]. Due to the presence of a time-like asymptotic boundary, Hawking radiation can be reflected back and equilibrium may be reached during the Hawking evaporation. There are some discussions on how the thermal equilibrium is reached for AdS black hole [11, 38], it would be an interesting topic to discuss the effect of Horndeski coupling on this process. Furthermore, as we have shown in this paper, the scalar charge introduced in Horndeski gravity is not an independent variable because the scalar field does not have its own dynamics. This physical property is not very important in thermodynamical first law but can significantly affect the black hole evaporation behavior as evaporation is a dynamic process. Thus it would be interesting to investigate the evaporating behavior for other scalar hair black holes where the scalar hair has its own dynamics, such as the case Ref.[33, 32, 34]. In these cases, the scalar charge will not follow Eq.(35) during evaporation but has its own evolution equation based on the dynamics. This certainly will offer more interesting phenomena.
@@ Acknowledgements
We are grateful to our group members for useful discussions. This work is supported by the National Natural Science Foundation of China (NSFC) under Grant Nos.12175105, 11575083, 11565017, and 12405066. Yu-Sen An is also supported by the Natural Science Foundation of Jiangsu Province under Grant No. BK20241376 and Ya-Peng Hu are also supported by the Top-notch Academic Programs Project of Jiangsu Higher Education Institutions (TAPP).

### v3

@@ 1 Introduction
Through indirect and direct observations by LIGO and the Event Horizon Telescope [1, 2, 3, 4], black holes, the most extraordinary objects in the universe, have received increasing attention. One of the most interesting things about black holes is that they behave like a thermodynamical system. In Ref.[5, 6], by using quantum field theory in curved space-time, Hawking showed that quantum effects allow black holes to radiate like black bodies which are now called Hawking radiation. From the Hawking radiation spectrum, one can calculate the power of the particle emission, investigate the evaporation behavior, and estimate the lifetime of black holes. After the discovery of Hawking radiation, Page gave numerical results for the lifetimes of uncharged and non-rotating black holes in asymptotically flat spacetime [7]. He found that there is a simple relation between a black hole’s lifetime and its initial mass $t\sim M_{0}^{3}$ [7], and then extended these discussions to rotating black hole case [8]. In Ref.[9], by combining the Hawking radiation and Schwinger effect in quantum electrodynamics, Hiscock and Weems generalized the discussions of black hole evaporation to the charged black hole case. They found that the presence of electric charge may greatly extend the lifetime of a black hole [9]. The study of the evaporation behavior of black holes has been a fruitful area up until now, it has been extended to many different spacetime backgrounds, such as de-Sitter spacetimes [10] and anti-de-Sitter spacetimes [11]. Due to the diversity of gravitational theories, there are also many discussions on black hole evaporation in modified gravity theories [12, 13, 14, 15, 16]. Moreover, black hole evaporation may also be of help in the search for dark matter and dark energy [17].
@@
Einstein’s general relativity where all the matter fields are minimally coupled to gravity has passed various direct or indirect tests. However, many extensions of general relativity have also been taken into account in order to explain the accelerated expansion of our universe [18, 19]. Particularly, a lot of work has been devoted to finding the most general non-minimal coupled scalar-tensor action whose equations of motion have at most second-order derivatives. In the 1970s, general classes of theories of this kind were constructed by Horndeski, which is now called Einstein-Horndeski gravity [20]. Although Horndeski gravity is a scalar-tensor theory, by gauging the shift symmetry of the scalar field, it is also equivalent to the Einstein-vector theory which is shown in Ref.[21]. Horndeski gravity has the coupling between the derivative of the scalar field and Einstein tensor. This feature makes Horndeski gravity theory useful in inflationary cosmology [22, 23] and late-time cosmology [24]. Besides investigations in the cosmological background, Horndeski gravity has also received much attention in static and spherically symmetric black hole systems. Various interesting topics have been discussed in Horndeski black hole background. To name a few, Ref.[25, 26, 27] discussed the Horndeski black hole thermodynamics, Ref.[28, 29, 30] discussed the holographic implication of Horndeski black hole, Ref.[31] focused on superradiation of Horndeski black hole and [32] discussed the black hole shadow in Horndeski background. Despite the above progress, black hole evaporation behavior which is an important part of black hole thermodynamics has not been investigated in Horndeski gravity background.
@@
Interestingly, from the equation of scalar field, a conserved current can be defined as
$J^{\mu}=(\alpha g^{\mu\nu}-\eta G^{\mu\nu})\nabla_{\nu}\phi,$ (4)
which can be used to prove the no-hair theorem as shown in Ref.[33]. In this work, we only consider the static and spherically symmetric black hole solution. So we can write the most general ansatz as
@@
in which the $\Phi(r)$ is the U(1) gauge field potential. By solving the Maxwell equation $\nabla_{\nu}F^{\nu\mu}=0$, the field potential $\Phi$ can be expressed as
$\Phi^{\prime}=\frac{4Q}{r^{2}}\sqrt{\frac{h}{f}},$ (6)
where $Q$ is an integration constant. By plugging the ansatz (5) into the above equations (2), there are three more independent equations which read
@@
$4(2rf^{\prime}+2(f-1)+2\Lambda r^{2})+\frac{32Q^{2}}{r^{2}}+2\alpha r^{2}f\phi^{\prime 2}$
$+2\eta(4rf\phi^{\prime\prime}+(3rf^{\prime}+(f+1))\phi^{\prime})f\phi^{\prime}$ $=0,$
$4(2rfh^{\prime}+2h(f-1)+2\Lambda r^{2}h)+\frac{32Q^{2}}{r^{2}}h-2\alpha r^{2}fh\phi^{\prime 2}$
@@
It should be pointed out that this equation is equivalent to $\alpha g^{rr}-\eta G^{rr}=0$, which is obviously related to Eq.(4). The underlying physical meaning will be discussed later.
For the case $\alpha\neq 0$ and $\eta\neq 0$, by analyzing the Riemann tensor near the spacetime infinity, the AdS radius can be defined as
$l^{-2}:=\frac{\alpha}{3\eta},$ (9)
where $\alpha/\eta$ is positive by the requirement of reality of the solution [34]. Thus for non-zero $\alpha$, the spherical black hole spacetime can only be the asymptotically AdS black hole. However in this paper, based on the observational consideration, we are more interested in asymptotically flat black hole solution. In order to get asymptotically flat black hole, we should set $\alpha=0$ which corresponds to the divergent AdS radius. So we will focus on the case where $\alpha=0$ and $\Lambda=0$ in the following. In this scenario, by directly solving the coupled differential equations Eq.(7) and (8), an asymptotically flat black hole solution can be obtained where [34, 27]
@@
$h(x)=1-\frac{2}{x}+\frac{4y^{2}}{x^{2}}-\frac{4y^{4}}{3x^{4}}.$ (13)
The charge parameter $y$ of this Horndeski black hole always satisfies the constraint Eq.(11). Under this constraint, we plot the $h(r)$ in Fig.1 to show the black hole metric function for different values of charge parameter $y$. It is easy to find that the metric function has only one line intersecting the $z=0$ plane which means that there is only one positive real root of Eq.(12). This means the charged Horndeski black hole we considered has only one event horizon. It should be stressed that this behavior with only one horizon is different from cases in Einstein gravity where the charged black holes and rotating black holes both have two horizons [35]. Interestingly, the rotating black hole in Horndeski gravity can also have both inner and outer horizons as shown in Ref.[36].
@@ 3 Thermodynamics of charged black holes in Einstein-Horndeski gravity
$\int_{c}\omega(\tilde{\phi},\delta\tilde{\phi},\mathcal{L}_{\xi}\tilde{\phi})=\delta\mathcal{H}_{\infty}-\delta\mathcal{H}_{+},$ (15)
@@
$\delta\mathcal{H}_{\infty}=\delta M-\Phi_{e}\delta Q,$ (16)
@@
$\delta\mathcal{H}_{+}=T\delta S+\Phi_{\phi}\delta Q_{\phi},$ (19)
where $\Phi_{\phi}$ is the corresponding conjugate quantity of $Q_{\phi}$. Due to this additional term, the thermodynamic first law is different from the RN case which reads
@@
$\Phi_{e}=\frac{4Q}{{r_{h}}}-\frac{8Q^{3}}{3r_{h}^{3}},$
where $\Phi_{e}$ is the electric potential difference between the infinity and the black hole event horizon $\Phi_{e}=\Phi|_{\infty}-\Phi|_{r_{h}}$, which is calculated from the third equation of the Eq.(10) by choosing the electric potential at the infinity as zero. The condition $\eta<0$ is required to ensure these thermodynamic variables are real. Moreover, unlike Einstein gravity, the entropy of the black hole in Horndeski gravity has correction which is
@@ 4 Evaporation behavior of charged black holes in Einstein-Horndeski gravity:
$(\alpha g^{rr}-\eta G^{rr})\phi^{\prime}\equiv 0,$ (24)
which means the $\phi$ field can not be solved by its equation of motion. This is very different from other cases where scalar hair is involved such as Ref.[37, 38, 39]. In these cases, the scalar field is solved by its equation of motion. It should be emphasized here that the Eq.(8) is important to escape from the constraints of no-hair theorem proposed in Ref.[33]. The proof of no hair theorem in Ref.[33] relies on the fact that the vanishing conserved current $J^{r}$ leads to vanishing scalar field $\phi$. However, by taking the condition Eq.(8) into account, the vanishing conserved current $J^{r}$ in Eq.(4) can not obtain the requirement of zero scalar field, which is crucial for the black hole to have non-trivial scalar hair as shown in Eq.(10). For more discussions on this point, the readers can consult Ref.[40, 41, 42, 43, 44, 45, 46].
@@
$\frac{dM}{dt}=T\frac{dS}{dt}+\Phi_{e}\frac{dQ}{dt}+\Phi_{\phi}\frac{dQ_{\phi}}{dt}.$ (28)
@@
The first term regarding thermal radiation follows Stefan-Boltzmann law
$\frac{\delta Q_{H}}{dt}=T\frac{dS}{dt}=-\sigma T^{4}.$ (29)
@@
$\frac{dr}{d\lambda}=\sqrt{\frac{f(r)}{b^{2}h(r)}-\frac{f(r)}{r^{2}}},$ (32)
where $b=L/\mathcal{E}$ is called the impact factor. A massless particle that can reach infinity must obey the equation
@@
$\frac{\mathrm{d}M}{\mathrm{d}t}=-\frac{\pi^{3}}{15}b_{c}^{2}T^{4}+\Phi_{e}\frac{\mathrm{d}Q}{\mathrm{d}t}+\Phi_{\phi}\frac{\mathrm{d}Q_{\phi}}{\mathrm{d}t},$ (37)
@@
$\frac{\mathrm{d}Q_{\phi}}{\mathrm{d}t}=\frac{\partial Q_{\phi}}{\partial M}\frac{\mathrm{d}M}{\mathrm{d}t}+\frac{\partial Q_{\phi}}{\partial Q}\frac{\mathrm{d}Q}{\mathrm{d}t}.$ (38)
@@
To understand the difference in evaporation behavior between RN and Horndeski black hole, we see that in combination with (40) and (42), the charge loss rate is closely related to the electric field strength outside the event horizon. At the early stage of evaporation, the horizon radius $r_{h}$ is large, thus $O(1/r^{4})$ term can be neglected in Eq.(41). Therefore, the electric field strength of Horndeski black holes outside the event horizon is the same as the RN case which is both $O(\frac{1}{r^{2}})$ Coulomb type. As the black hole evaporates, $r_{h}$ becomes smaller and the $O(\frac{1}{r^{4}})$ term of Horndeski electric field strength (41) becomes more and more important. As can be seen from Fig.2, at the late stage of evaporation, the exterior electric field strength of the Horndeski black hole is vastly different from the RN case. The electric field strength outside Horndeski black hole first increases and then decreases instead of being a monotonically decreasing function. This difference in the electric field strength is the crucial reason for why the evaporation behavior of Horndeski black holes deviate from the RN black holes.
@@ 5 Conclusion
For future directions, firstly, we only consider the asymptotically flat Horndeski black hole in this work. For Horndeski gravity, there are also black hole solutions in asymptotically AdS spacetime [27]. Due to the presence of a time-like asymptotic boundary, Hawking radiation can be reflected back and equilibrium may be reached during the Hawking evaporation. There are some discussions on how the thermal equilibrium is reached for AdS black hole [11, 50], it would be an interesting topic to discuss the effect of Horndeski coupling on this process. Moreover, for simplicity, we only consider the black hole evaporation in a special truncation of Horndeski gravity theory. The generalization of our analysis to more general scalar-tensor theories are still interesting to be pursued. Finally, as we have shown in this paper, the scalar charge introduced in Horndeski gravity is not an independent variable because the scalar field does not have its own dynamics. This physical property is not very important in thermodynamical first law but can significantly affect the black hole evaporation behavior as evaporation is a dynamic process. Thus it would be interesting to investigate the evaporating behavior for other scalar hair black holes where the scalar hair has its own dynamics, such as the case Ref.[38, 37, 39]. In these cases, the scalar charge will not follow Eq.(38) during evaporation but has its own evolution equation based on the dynamics. This certainly will offer more interesting phenomena.
@@ Acknowledgements
We are grateful to our group members for useful discussions. This work is supported by the National Natural Science Foundation of China (NSFC) under Grant Nos.12175105, 12405066, 11575083, and 11565017. Yu-Sen An is also supported by the Natural Science Foundation of Jiangsu Province under Grant No.BK20241376 and Fundamental Research Funds for the Central Universities. Ya-Peng Hu is also supported by the Top-notch Academic Programs Project of Jiangsu Higher Education Institutions (TAPP).